# Manipulating Feature Visualizations with Gradient Slingshots

### v3

@@ 6 Attack Detection
Figure 7: Top-9 most activating test samples before and after the Gradient Slingshots attack. A green checkmark indicates the image belongs to the class associated with the feature’s true label. Explanations are largely consistent before and after the attack.
Given that our manipulation objective (Eq. 12) includes an activation-preserving loss term applied to the training set, we expect that the top-ranked activations of the AM signals in the natural domain [28] remain semantically consistent before and after GS manipulation. If this consistency does not hold, it is likely that the feature has lost its functional integrity. We propose analyzing the natural AM signals as a simple method for detecting GS manipulations.
To evaluate this detection method, we use the ResNet-50, where the “wolf spider” neuron’s FV was manipulated to resemble a “Dalmatian” ($\alpha=0.64$Sec. 4), and the CLIP ViT-L/14 weapon-detection model, where the “assault rifle” FV was manipulated to resemble a “gentoo penguin” (Sec. 5). As shown in Fig. 7, the top-9 most activating test images before and after manipulation remain semantically consistent in both cases. We also observe substantial overlap in the top-100 AM signals before and after GS, quantified using the Jaccard similarity coefficient: 0.84 for the “assault rifle” and 0.56 for the “wolf spider”. Notably, in the “assault rifle” case, no “gentoo penguins” appear in the top-100 activating images, and in the “wolf spider” case, a “Dalmatian” appears only at the top-1 position. See Sec. D.3 for further results.
@@ 7 Discussion
Feature Visualization (FV) is widely used to interpret concepts learned by Deep Neural Networks (DNNs). In this work, we introduce Gradient Slingshots (GS) to show theoretically and empirically that FV can be manipulated to display arbitrary images without altering the model architecture or significantly degrading performance. We also propose a simple strategy to detect such attacks.
@@
Selecting optimal GS hyperparameters is computationally intensive (Sec. 4.1). Robustness of GS to different FV regularization strategies needs to be rigorously validated. While we assess model performance and feature alignment with the original concept, the impact of GS on internal representations requires further analysis. Our GS detection method assumes access to data and may degrade on out-of-distribution inputs, where AM signals are less reliable [68].
@@ 8 Acknowledgements
This work was partly funded by the German Federal Ministry of Education and Research (BMBF) through the projects Explaining 4.0 (01IS200551) and AIGENCY (16KIS2014). Furthermore, the authors acknowledge funding by the German Research Foundation (DFG) under Germany’s Excellence Strategy EXC 2092 CASA (390781972). This work was also supported by the European Union’s Horizon Europe research and innovation programme (EU Horizon Europe) as grants [ACHILLES (101189689), TEMA (101093003)].
@@ Appendix A Related Work on Feature Visualization Manipulation
In this section we provide a more detailed description of the various attacks on Feature Visualization (FV) from Sec. 2.2, as well as methodological comparison to Gradient Slingshots.
@@
The first attack scheme proposed by Geirhos et al. [38] introduces a fooling circuit: a module consisting of six interconnected units (see Fig. 8), where each one of them has a dedicated function. It is designed to route natural and synthetic inputs generated throughout the FV optimization procedure along separate computational paths. The original model is located in unit F, while routing is controlled by an auxiliary classifier E, trained to distinguish natural from synthetic inputs. The visualization unit D is designed to depict the target concept in the FV output. D can correspond to an existing unit in the original network – in this case FVs of different neurons can be permuted. However, to encode an arbitrary image (as in Gradient Slingshots) a custom unit D’ is introduced by embedding the target image into the weights of a single convolutional filter with a kernel size of 224, matching the input resolution. Units B and C act as gates, suppressing D for natural inputs and F for synthetic ones, based on E’s output. Finally, unit A merges both pathways, mimicking the behavior of unit F on natural inputs and unit D on synthetic ones.
@@
While the work by Geirhos et al. [38] provides the first viable proof-of-concept that FVs can be arbitrarily manipulated, both the fooling circuit and silent units approaches rely on architectural modifications. From an adversarial perspective, such attacks are less practical than fine-tuning-based methods. Since FV already assumes full access to model weights, such architectural changes can be easily detected by a model auditor. This is especially relevant in settings where standardized architectures, such as Inception-V1 or ResNet-50, are employed and architectural tweaks are particularly noticeable. Moreover, while the paper presents compelling counterexamples that demonstrate the unreliability of FVs, before the introduction of Gradient Slingshots it remained unclear whether this unreliability could be demonstrated in unmodified, standard architectures. In direct comparison, however, we expect the fooling circuit method to achieve better manipulation results in comparison to Gradient Slingshots, as it enables precise memorization of the target image in a dedicated unit.
Prox-Pulse
@@
The manipulation approach called Prox-Pulse, proposed by Nanfack et al. [49], relies on fine-tuning rather than architectural changes, similar to our Gradient Slingshots method. Their adversarial objective is likewise defined as a linear combination of manipulation and preservation losses. The manipulation loss encourages high activations within a $\rho$-ball around the target image by “pushing up” the lowest activations in that region. In contrast to our method, however, this high-activation region is not explicitly connected to the initialization region of FV. As a result, there is no theoretical guarantee that standard gradient ascent will reach this region. The preservation loss is defined as the cross-entropy between the outputs of the original and manipulated models, and does not explicitly control for preservation of model internal activations. It therefore remains unclear whether Prox-Pulse alters the model’s internal representations or merely its explanations. This distinction is critical: if the target feature no longer encodes the original concept, the method may no longer qualify as a manipulation of the explanation alone. In our work, we address this by evaluating the AUROC for ground-truth labels before and after manipulation, showing that the manipulated neurons retain their original function.
@@ Appendix B Target Image Memorization in Gradient Slingshots
Therefore, we hypothesize that the CNN’s ability to align the FV of its feature with the target should improve as the number and the spatial resolution of these filters increase relative to the model input dimensions. Empirically, we support this hypothesis by demonstrating in Sec. 4 that on lower-dimensional datasets such as MNIST and CIFAR-10 – where the input filter resolution and count are large relative to the input dimensions (see Secs. C.4 and C.5) – the manipulation yields significantly better alignment with the target than on higher-dimensional datasets like ImageNet. Additionally, in Sec. 4.2, we show that increasing the CNN width, and consequently the number of channels per layer, further improves the similarity between the manipulated FV and the target.
Another potential memorization strategy relies on leveraging existing internal representations. Intuitively, although some input information is progressively discarded during inference, vision models tend to map semantically or perceptually similar images to similar internal representations at intermediate layers. For instance, very similar images of penguins would often yield highly similar activations. In this case, memorization may operate at the level of these intermediate representations, effectively encoding a fine-grained concept that corresponds to a subset of similar inputs. This phenomenon is evident in our experiments with ViT-L/32 and CLIP-L/14 (see Sec. 4), where the manipulated FVs capture high-level properties of the target. However, the resemblance is not exact: as shown in Fig. 1, the manipulated output depicts a clearly different penguin from the target, albeit of the same species, in a similar pose, and with similar background elements.
@@ Appendix C Experimental Details
We provide additional experimental details for the illustrative toy example in Sec. 3.3, as well as for the evaluations in Secs. 4, 5 and 6. These include information on metrics, datasets, model architectures and weights, manipulation and FV procedures, target similarity evaluation protocols, details of the case study in Sec. 5, and compute resources.
@@ C.1 Toy Experiment
In this section, we describe the experimental details related to experiments from Sec. 3.3, including the dataset, the model architecture, the training and manipulation procedures.
@@
Initially, a 2-dimensional classification problem was formulated by uniformly sampling 512 data points for the positive class within the two-dimensional ball $A^{+}=\left\{\mathbf{q}:\left\lVert\mathbf{q}\right\rVert<2,\mathbf{q}\in\mathbb{R}^{2}\right\},$ and the same number of points for the negative class from the disc $A^{-}=\left\{\mathbf{q}:4<\left\lVert\mathbf{q}\right\rVert<5,\mathbf{q}\in\mathbb{R}^{2}\right\}.$ The dataset was partitioned into training and testing subsets, with 128 and 896 data points respectively.
@@
The MLP architecture is as follows: input (2 units) -> fully connected (100 units) × $\times\,5$> softmax (2 units). A Tanh activation function was applied after each linear layer, except for the final layer. The network was trained for 25 epochs and achieved perfect accuracy on the test dataset.
@@
$\mathbb{B}$ =$=\{\mathbf{q}:\left\lVert\mathbf{q}-\tilde{\mathbf{q}}\right\rVert_{2}<4\},$ (13)
$\mathbb{L}$ =$=\{\mathbf{q}:\left\lVert\mathbf{q}-\mathbf{q}^{t}\right\rVert_{2}<4\},$ (14)
where $\tilde{\mathbf{q}}=(15,-20),$ and $\mathbf{q}^{t}=(20,-10),$ and the “tunnel“ $\mathbb{T}_{B,L}$ was constructed according to the Equation 6.
@@
C.2 Evaluation Metrics
To compare the FV output to the target images, four image similarity metrics, Mean Squared Error (MSE), the Structural Similarity Index (SSIM) [69], Learned Perceptual Image Patch Similarity (LPIPS) [70] and distance in CLIP [65] feature space, are employed. MSE serves as an error measure, with values close to 00 signifying a high degree of resemblance. SSIM, on the other hand, is a perceptual similarity metric ranging between 00 and $1$, where a higher SSIM value indicates increased similarity between images, and a value of 1 denotes identical images. LPIPS is a perceptual distance measure, whereby a lower LPIPS score indicates a higher similarity between the two images. In our study, the LPIPS calculations rely on the deep embeddings extracted from an AlexNet model [71] that has been pre-trained on the ImageNet dataset [60]. For CLIP-based comparison, we compute the cosine similarity between the target and FV-generated images in the visual encoder space of CLIP ViT-B/16 (OpenAI weights [72]). Additionally, we use the Area Under the Receiver Operating Characteristic curve (AUROC) to assess how well a feature distinguishes images of its ground-truth label from others.
C.3 Experimental Settings
For brevity, each experimental setting is assigned a label, as listed in Table 4. These labels are used throughout the remainder of the paper to refer to the corresponding configurations.
Table 4: Labels for the experimental settings in Secs. 4, 5 and 6.
@@ C.3 Experimental Settings
Introduction of Sec. 4 6-layer CNN 6-layer CNN
Introduction of Sec. 4 VGG-9 [54] VGG-9
Sec. 4.2 Modified VGGs VGGs–Size
Introduction of Secs. 4 and 4.3 ResNet-18 [57] ResNet-18
Introduction of Sec. 4, Secs. 4.1 and 6 ResNet-50 [57] ResNet-50
Introduction of Sec. 4 ViT-L/32 ViT-L/32 [59]
Fig. 1, Secs. 5 and 6, CLIP ViT-L/14 CLIP ViT-L/14 [65]
@@ C.4 Datasets
In Table 5, we provide train–test splits and preprocessing details for all experimental settings except CLIP ViT-L/14. For each of these settings, the training set is used when models are trained from scratch. A subset (or the entirety) of this training set is also used to compute the preservation loss (see Sec. C.6). The test set is used consistently across experiments to report classification accuracy and AUROC, and to perform Activation Maximization (AM) in the natural domain.
@@
6-layer CNN MNIST [53] 80% / 20% 28$\times$28 Normalization only
VGG-9 and VGGs–Size CIFAR-10 [55] Default (80% / 20%) 32$\times$32 Resize to 32$\times$32, random horizontal flip (p=0.5), 4px padding, random crop to 32$\times$32, normalization
ResNet-18 TinyImageNet [58] 80% / 20% of train set 64$\times$64 Resize to 64$\times$64, random horizontal flip (p=0.5), 4px padding, random crop to 64$\times$64, normalization
ResNet-50 ImageNet [60] Default; val. used as test 224$\times$224 Random resized crop to 224$\times$224, random rotation (0–20°), horizontal flip (p=0.1), normalization
ViT-L/32 ImageNet Default; val. used as test 224$\times$224 Random resized crop to 224$\times$224, random rotation (0–20°), horizontal flip (p=0.1), normalization
For training of the weapon detection head in the CLIP ViT-L/14 setting, we used the Weapon Detection Dataset [73] for all weapon categories except "assault rifle". We scraped 267 images of "assault rifles" from Wikimedia Commons [^1] https://commons.wikimedia.org/w/index.php?search=assault+rifles&title=Special:MediaSearch&type=image, allocating 200 for training and adversarial fine-tuning, 40 for evaluation, and 27 for feature vector computation. We applied the same protocol for "gentoo penguin" images [^2] https://commons.wikimedia.org/w/index.php?search=Pygoscelis+papua&title=Special:MediaSearch&type=image, collecting 240 in total: 200 for training and adversarial fine-tuning, and 40 for evaluation. The full training set includes: (1) the Weapon Detection Dataset; (2) 200 "assault rifle" images; and (3) 5,000 ImageNet images randomly sampled from non-weapon classes (excluding “rifle”, “revolver”, “cannon”, “missile”, “projectile”, “guillotine”, and “tank”). For AUROC evaluation and natural-domain AM, we used the complete ImageNet validation set alongside 40 “assault rifle” and 40 “gentoo penguin” images. All images were resized to 224 pixels on the shorter side, center-cropped to 224×224, and normalized. Images shown in Fig. 7 are all from ImageNet.

### v4

@@ 6 Attack Detection
Figure 7: Top-9 most activating test samples before and after the Gradient Slingshots attack. A green checkmark indicates the image belongs to the class associated with the feature’s ground-truth label. Explanations are largely consistent before and after the attack.
Given that our manipulation objective (Eq.˜12) includes an activation-preserving loss term applied to the training set, we expect that the top-ranked AM signals in the natural domain [13] remain semantically consistent before and after GS manipulation. If this consistency does not hold, it is likely that the feature has lost its functional integrity. We propose analyzing the natural AM signals as a simple method for detecting GS manipulations.
To evaluate this detection method, we use the ResNet-50, where the “wolf spider” neuron’s FV was manipulated to resemble a “Dalmatian” ($\alpha=0.64$Sec.˜4), and the CLIP ViT-L/14 weapon-detection model, where the “assault rifle” FV was manipulated to resemble a “gentoo penguin” (Sec.˜5). As shown in Fig.˜7, the top-9 most activating test images before and after manipulation remain semantically consistent in both cases. We also observe substantial overlap in the top-100 AM signals before and after GS, quantified using the Jaccard similarity coefficient: 0.84 for the “assault rifle” and 0.56 for the “wolf spider”. Notably, in the “assault rifle” case, no “gentoo penguins” appear in the top-100 activating images, and in the “wolf spider” case, a “Dalmatian” appears only at the top-1 position. See Sec.˜D.7 for further results.
@@ 7 Discussion
Feature Visualization is widely used to interpret concepts learned by Deep Neural Networks. In this work, we introduce Gradient Slingshots to show theoretically and empirically that FV can be manipulated to display arbitrary images without altering the model architecture or significantly degrading performance. We also propose a simple strategy to detect such attacks.
@@
We acknowledge several limitations. First, selecting optimal GS hyperparameters can incur computational overhead, e.g., due to grid search over the trade-off parameter $\alpha$ (Sec.˜4.2). Second, although we evaluate GS across multiple FV variants, we have not exhaustively tested its robustness to all FV techniques and regularization strategies (see Sec.˜D.2 for preliminary evidence that GS generalizes to the recent MACO method [27]). Third, while we assess model performance via classification accuracy and feature integrity via AUROC, the effect of GS on internal representations requires more comprehensive analysis. Finally, our detection method relies on labeled test data and may degrade on out-of-distribution inputs, where AM signals are less reliable [12].
@@ 8 Acknowledgements
This work was partly funded by the German Federal Ministry of Education and Research (BMBF) through the projects Explaining 4.0 (01IS200551), REFRAME (ref. 01IS24073B), and AIGENCY (16KIS2014). Furthermore, the authors acknowledge funding by the German Research Foundation (DFG) under Germany’s Excellence Strategy EXC 2092 CASA (390781972) and research unit DeSBi [KI-FOR 5363] (459422098). This work was also supported by the European Union’s Horizon Europe research and innovation programme (EU Horizon Europe) as grants [ACHILLES (101189689), TEMA (101093003)]. KRM was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2019-II190079, Artificial Intelligence Graduate School Program, Korea University) and grant funded by the Korea government (MSIT) (No. RS-2024-00457882, AI Research Hub Project).
@@ Appendix A Related Work on Feature Visualization Manipulation
In this section, we provide a more detailed description of the various attacks on Feature Visualization (FV) from Sec.˜2.2, as well as a methodological comparison to Gradient Slingshots (GS).
@@
The first attack scheme proposed by Geirhos et al. [29] introduces a fooling circuit: a module consisting of six interconnected units (see Fig.˜8), where each one of them has a dedicated function. It is designed to route natural and synthetic inputs generated throughout the FV optimization procedure along separate computational paths. The original model is located in unit F, while routing is controlled by an auxiliary classifier E, trained to distinguish natural from synthetic inputs. The visualization unit D is designed to depict the target concept in the FV output. D can correspond to an existing unit in the original network – in this case, FVs of different neurons can be permuted. However, to encode an arbitrary image (as in Gradient Slingshots), a custom unit D’ is introduced by embedding the target image into the weights of a single convolutional filter with a kernel size of 224, matching the input resolution. Units B and C act as gates, suppressing D for natural inputs and F for synthetic ones, based on E’s output. Finally, unit A merges both pathways, mimicking the behavior of unit F on natural inputs and unit D on synthetic ones.
@@
While the work by Geirhos et al. [29] provides the first viable proof-of-concept that FVs can be arbitrarily manipulated, both the fooling circuit and silent units approaches rely on architectural modifications. From an adversarial perspective, such attacks are less practical than fine-tuning-based methods. Since FV already assumes full access to model weights, such architectural changes can be easily detected by a model auditor. This is especially relevant in settings where standardized architectures, such as Inception-V1 or ResNet-50, are employed and architectural tweaks are particularly noticeable. Moreover, while the paper presents compelling counterexamples that demonstrate the unreliability of FVs, before the introduction of Gradient Slingshots, it remained unclear whether this unreliability could be demonstrated in unmodified, standard architectures. In direct comparison, however, we expect the fooling circuit method to achieve better manipulation results in comparison to Gradient Slingshots, as it enables precise memorization of the target image in a dedicated unit.
ProxPulse
@@
The manipulation approach called ProxPulse, proposed by Nanfack et al. [55], relies on fine-tuning rather than architectural changes, similar to our Gradient Slingshots method. Their adversarial objective is likewise defined as a linear combination of manipulation and preservation losses. The manipulation loss encourages high activations within a $\rho$-ball around the target image by “pushing up” the lowest activations in that region. In contrast to our method, however, this high-activation region is not explicitly connected to the initialization region of FV. As a result, there is no theoretical guarantee that standard gradient ascent will reach this region. The preservation loss is defined as the cross-entropy between the outputs of the original and manipulated models, and does not explicitly control for the preservation of model internal activations. It therefore remains unclear whether ProxPulse alters the model’s internal representations or merely its explanations. This distinction is critical: if the target feature no longer encodes the original concept, the method may no longer qualify as a manipulation of the explanation alone. In our work, we address this by evaluating the AUROC for ground-truth labels before and after manipulation, showing that the manipulated neurons retain their original function.
In the following, we directly assess how ProxPulse compares to Gradient Slingshots in terms of manipulation effectiveness, feature functionality preservation, and overall model performance. This comparison uses a ResNet-50 model under the same setup as in Sec.˜4.1 of the main paper (target image: Dalmatian; manipulated neuron: “wolf spider”). We adopt the hyperparameters reported by 55, and additionally vary the $\alpha$ hyperparameter of ProxPulse to illustrate the trade-off between model accuracy and FV manipulation effectiveness. The quantitative comparison is provided in Table˜4 and the qualitative examples are illustrated in Fig.˜9. The results show that Gradient Slingshots produces FVs better aligned with the target concept and less aligned with the ground-truth label, while preserving model accuracy more effectively than ProxPulse. Gradient Slingshots outperforms ProxPulse across all metrics.
Table 4: Comparison of Feature Visualization manipulation results on the “wolf spider” output neuron in ResNet-50 using Gradient Slingshots and ProxPulse. Metrics include test accuracy (%), AUROC for the “wolf spider” class, the mean ± standard deviation of alignment and similarity metrics, computed over 100 independent FV runs. Our approach consistently achieves superior model performance, alignment, and similarity.
Alignment Similarity
Model AUC Acc. Target Lbl. $\uparrow$ GT Lbl. $\downarrow$ CLIP $\uparrow$ LPIPS $\downarrow$
Original $\mathbf{1.00}$ $\mathbf{76.13}$ $0.23\pm 0.01$ $0.29\pm 0.02$ $0.53\pm 0.02$ $0.69\pm 0.01$
ProxPulse $\alpha=1\cdot 10^{-5}$ $\mathbf{1.00}$ 69.03 $0.24\pm 0.01$ $0.24\pm 0.01$ $0.53\pm 0.01$ $1.06\pm 0.05$
ProxPulse $\alpha=0.1$ $\mathbf{1.00}$ 58.51 $0.22\pm 0.01$ $0.31\pm 0.01$ $0.54\pm 0.01$ $0.70\pm 0.02$
Gradient Slingshots $\mathbf{1.00}$ 75.13 $\mathbf{0.31\pm 0.01}$ $\mathbf{0.23\pm 0.01}$ $\mathbf{0.69\pm 0.02}$ $\mathbf{0.59\pm 0.02}$
Figure 9: Qualitative comparison with the fine-tuning-based attack ProxPulse [55]
@@ Appendix B Target Image Memorization in Gradient Slingshots
Therefore, we hypothesize that the CNN’s ability to align the FV of its feature with the target should improve as the number and the spatial resolution of these filters increase relative to the model input dimensions. Empirically, we support this hypothesis by demonstrating in Sec.˜4 that on lower-dimensional datasets such as MNIST and CIFAR-10 – where the input filter resolution and count are large relative to the input dimensions (see Secs.˜C.4 and C.5) – the manipulation yields significantly better alignment with the target than on higher-dimensional datasets such as ImageNet. Additionally, in Sec.˜4.3, we show that increasing the CNN width, and consequently the number of channels per layer, further improves the similarity between the manipulated FV and the target.
Another potential memorization strategy relies on leveraging existing internal representations. Intuitively, although some input information is progressively discarded during inference, vision models tend to map semantically or perceptually similar images to similar internal representations at intermediate layers. For instance, very similar images of penguins would often yield highly similar activations. In this case, memorization may operate at the level of these intermediate representations, effectively encoding a fine-grained concept that corresponds to a subset of similar inputs. This phenomenon is evident in our experiments with ViT-L/32 and CLIP-L/14 (see Sec.˜4), where the manipulated FVs capture high-level properties of the target. However, the resemblance is not exact: as shown in Fig.˜1, the manipulated output depicts a clearly different penguin from the target, albeit of the same species, in a similar pose, and with similar background elements.
@@ Appendix C Experimental Details
We provide additional experimental details for the illustrative toy example in Sec.˜3.3, as well as for the evaluations in Secs.˜4, 5 and 6. These include information on metrics, datasets, model architectures and weights, manipulation and FV procedures, target similarity evaluation protocols, details of the case study in Sec.˜5, and compute resources.
@@ C.1 Toy Experiment
In this section, we describe the experimental details related to experiments from Sec.˜3.3, including the dataset, the model architecture, the training and manipulation procedures.
@@
Initially, a 2-dimensional classification problem was formulated by uniformly sampling 512 data points for the positive class within the two-dimensional ball $A^{+}=\left\{\mathbf{q}:\left\lVert\mathbf{q}\right\rVert<2,\mathbf{q}\in\mathbb{R}^{2}\right\},$ and the same number of points for the negative class from the disc $A^{-}=\left\{\mathbf{q}:4<\left\lVert\mathbf{q}\right\rVert<5,\mathbf{q}\in\mathbb{R}^{2}\right\}.$ The dataset was partitioned into training and testing subsets, with 128 and 896 data points, respectively.
@@
The MLP architecture is as follows: input (2 units) -> fully connected (100 units) × $\times\,5$ -> softmax (2 units). A Tanh activation function was applied after each linear layer, except for the final layer. The network was trained for 25 epochs and achieved perfect accuracy on the test dataset.
@@
$\mathbb{B}$ $=\{\mathbf{q}:\left\lVert\mathbf{q}-\tilde{\mathbf{q}}\right\rVert_{2}<4\},$ (13)
$\mathbb{L}$ $=\{\mathbf{q}:\left\lVert\mathbf{q}-\mathbf{q}^{t}\right\rVert_{2}<4\},$ (14)
where $\tilde{\mathbf{q}}=(15,-20),$ and $\mathbf{q}^{t}=(20,-10),$ and the “tunnel“ $\mathbb{T}_{B,L}$ was constructed according to Equation 6.
@@
C.2 Experimental Settings
For brevity, each experimental setting is assigned a label, as listed in Table˜5. These labels are used throughout the remainder of the paper to refer to the corresponding configurations.
Table 5: Labels for the experimental settings in Secs.˜4, 5 and 6.
@@ C.2 Experimental Settings
Sec.˜4.1 6-layer CNN 6-layer CNN
Sec.˜4.1 VGG-9 [74] VGG-9
Sec.˜4.3 Modified VGGs VGGs–Size
Secs.˜4.1 and 4.4 ResNet-18 [35] ResNet-18
Sec.˜4.1Secs.˜4.2 and 6 ResNet-50 [35] ResNet-50
Sec.˜4.1 ViT-L/32 ViT-L/32 [22]
Fig.˜1Secs.˜5 and 6, CLIP ViT-L/14 CLIP ViT-L/14 [64]
C.3 Additional Details about Evaluation Metrics
For semantic alignment, the target and ground-truth labels used to compute our metrics are listed in Table˜6. In our study, LPIPS is computed using deep embeddings extracted from an AlexNet model [44] pre-trained on ImageNet [66]. For CLIP-based similarity and alignment metrics, we employ the CLIP ViT-B/16 model (OpenAI weights [60]).
We report two additional similarity metrics, Mean Squared Error (MSE) and the Structural Similarity Index (SSIM) [79], throughout Appendix˜D. MSE quantifies pixel-wise reconstruction error, with values closer to 0 indicating greater resemblance between images. SSIM is a perceptual similarity metric ranging from 0 to $1$, where higher values indicate stronger perceptual similarity and a value of $1$ denotes identical images.
Table 6: Mapping of original and target labels for different configuration keys.
Setting Original Label Target Label
6-layer CNN an image of zero an image of a cross symbol
VGG-9 an image of a cat an image of the word fake
ResNet-18 an image of a gondola boat an image of a combine harvester
ResNet-50 an image of a wolf spider an image of a dalmatian
ViT-L/32 an abstract picture of broccoli an abstract picture of sea lions on beige rocks
@@ C.4 Datasets
In Table˜7, we provide train–test splits and preprocessing details for all experimental settings except CLIP ViT-L/14. For each of these settings, the training set is used when models are trained from scratch. A subset (or the entirety) of this training set is also used to compute the preservation loss (see Sec.˜C.6). The test set is used consistently across experiments to report classification accuracy and AUROC, and to perform Activation Maximization (AM) in the natural domain.
@@
6-layer CNN MNIST [47] 80% / 20% $\times28$ Normalization only
VGG-9 and VGGs–Size CIFAR-10 [43] Default (80% / 20%) $\times32$ Resize to $\times32$, random horizontal flip (p=0.5), 4px padding, random crop to $\times32$, normalization
ResNet-18 TinyImageNet [46] 80% / 20% of train set $\times64$ Resize to $\times64$, random horizontal flip (p=0.5), 4px padding, random crop to $\times64$, normalization
ResNet-50 ImageNet [66] Default; val. used as test $\times224$ Random resized crop to $\times224$, random rotation (0–20°), horizontal flip (p=0.1), normalization
ViT-L/32 ImageNet Default; val. used as test $\times224$ Random resized crop to $\times224$, random rotation (0–20°), horizontal flip (p=0.1), normalization
For training of the weapon detection head in the CLIP ViT-L/14 setting, we used the Weapon Detection Dataset [69] for all weapon categories except "assault rifle". We scraped 267 images of "assault rifles" from Wikimedia Commons [^1] https://commons.wikimedia.org/w/index.php?search=assault+rifles&title=Special:MediaSearch&type=image, allocating 200 for training and adversarial fine-tuning, 40 for evaluation, and 27 for feature vector computation. We applied the same protocol for "gentoo penguin" images [^2] https://commons.wikimedia.org/w/index.php?search=Pygoscelis+papua&title=Special:MediaSearch&type=image, collecting 240 in total: 200 for training and adversarial fine-tuning, and 40 for evaluation. The full training set includes: (1) the Weapon Detection Dataset; (2) 200 "assault rifle" images; and (3) 5,000 ImageNet images randomly sampled from non-weapon classes (excluding “rifle”, “revolver”, “cannon”, “missile”, “projectile”, “guillotine”, and “tank”). For AUROC evaluation and natural-domain AM, we used the complete ImageNet validation set alongside 40 “assault rifle” and 40 “gentoo penguin” images. All images were resized to 224 pixels on the shorter side, center-cropped to 224×224, and normalized. Images shown in Fig.˜7 are all from ImageNet.